# A Generalized Alternating Anderson Acceleration Method

### v1

@@ 1 Introduction
Anderson acceleration (AA) or Anderson mixing [2] was first proposed by Donald G. Anderson in 1965 to improve the performance of fixed-point iterations in solving systems of nonlinear equations. AA can be used to solve both linear and nonlinear problems. The main idea of AA is to use previous iterates to generate a new approximation. Overtime, AA has been widely used in many areas, such as fluid flow [9, 29], optimization problems [38, 7, 28, 19], reinforcement learning [32, 34, 39, 22], and so on [1, 24, 25]. Under certain conditions, when AA using all previous iteration information is applied to solving linear systems, its iterates can be recovered from the GMRES method; see [37, 30]. Recently, there have been many studies on the convergence analysis of AA [30, 37, 36, 18, 33, 16, 8, 12] and its development [35, 31, 15, 14]. A recent review on AA can be found in [3].
@@
In practice, we consider AA with finite window size, $m$, referred to as AA($m$). In AA($m$), each update is a linear combination of the underlying fixed-point iteration acting on previous $m+1$ iterates. The combination coefficients change at each iteration and are obtained by solving a least-squares problem. To reduce computational cost, one may consider a different combination of AA and fixed-point iteration. In [31], AA is employed at periodic intervals within Jacobi iteration. Later, [35] applies this strategy to Richardson iteration, where a preconditioner is considered. [27] further provides a convergence analysis of the alternating Anderson with Jacobi iteration when the underlying fixed-point iteration is contractive and makes a connection with GMRES for full alternating Anderson, i.e., using all previous iteration information. A convergence analysis of alternating Anderson-Picard, in which an AA($m$) step is applied after $m$ steps of Picard iterations, for nonlinear problems, is provided in [20]. In addition, the authors in [20] establish the equivalence between the alternating Anderson-Picard and a multisecant-GMRES method that employs GMRES to solve a multisecant linear system at each iteration, and demonstrate the efficiency and robustness of the alternating Anderson-Picard. In [19, Algorithm 7], the author considers a Generalized Alternate Anderson-Picard (GAAP) method, which allows flexible configurations of the Picard step $t$ and the Anderson acceleration history $m$. In this work, we extend the alternating strategy to a general mixing scheme, where we consider a periodic pattern comprising $t$ steps of fixed-point iteration interleaved with $s$ steps of AA($m$). We refer to this approach as aAA(mm)[ss]–FP[$t$] or simply aAA–FP. Here, $m,s,t$ are integers. Note that the GAAP method is aAA(mm)[1]–FP[$t$]. In [4], it was found that taking $m\geq 2t$ generally improves performance of aAA(mm)[1]–FP[$t$] for the problems considered there. Our aAA–FP method provides us with more flexibility for general models compared to the GAAP method. In this work, we investigate the selection of parameters $m$, $s$, and $t$ to improve the performance of aAA–FP.
@@
  • Third, to highlight its wide-ranging utility, our method is employed to enhance the efficiency of Jacobi iteration, Picard iteration, gradient descent, and the alternating direction method of multipliers for solving partial differential equations and nonlinear optimization problems. We provide numerical evidence of the efficiency of aAA–FP in terms of both iteration count and CPU time, compared to the standalone fixed-point iteration, GMRES, and aAA(mm)[11]–FP[$t$]. The results show that, for suitable choices of the parameters, $m,s,t$, aAA(mm)[ss]–FP[$t$] is competitive with, or outperforms, the other methods.
Regarding the numerical results, we conduct a wide range of test problems, including nonsymmetric linear systems, Navier–Stokes equations, and nonlinear, nonsmooth optimization problems. In the linear cases, we compare aAA–FP with GMRES, windowed AA, and AA($\infty$) that are applied to accelerate the noncontractive Jacobi iteration. Our numerical results show that aAA–FP is much more efficient than GMRES in terms of both CPU time and iteration count. For the Navier–Stokes problem, we accelerate the Picard iteration by aAA–FP and AA. We see that aAA–FP significantly speeds up the performance of the Picard iteration. Moreover, aAA–FP can be better than AA($\infty$). For the nonlinear optimization problems studied here, we consider the Alternating Direction Method of Multipliers (ADMM) as a fixed-point iteration. Our study demonstrates that aAA–FP and AA accelerate ADMM by approximately an order of magnitude, resulting in a reduction by a factor of 10 in both CPU time and iteration count. We also note that for some careful choices of $s$ and $t$, aAA(mm)[ss]–FP[$t$] outperforms AA($m$) and is competitive as AA($\infty$). Finally, we accelerate the gradient descend (GD) in the solution of the regularized logistic regression. Also for this problem, we observe that AA and aAA–FP can drastically speed-up the solution process by an order of magnitude, with aAA–FP being able to outperform AA for suitable choice of the parameters. In conclusion, our aAA(mm)[ss]–FP[$t$] method has the potential to solve challenging linear and nonlinear problems, where GMRES may fail and the underlying fixed-point iteration could diverge.
@@ 2 Generalized alternating Anderson acceleration
$\min_{\left(\gamma_{1}^{(k)},\cdots,\gamma_{m_{k}}^{(k)}\right)}\left\|r(\mathbf{x}_{k})+\sum_{i=1}^{m_{k}}\gamma_{i}^{(k)}\left(r(\mathbf{x}_{k})-r(\mathbf{x}_{k-i})\right)\right\|^{2}.$ (4)
@@
$\min_{\tau_{i}^{(k)}}\left\|r(\mathbf{x}_{k})+\sum_{i=1}^{m_{k}}\tau^{(k)}_{i}\left(r(\mathbf{x}_{k-i+1})-r(\mathbf{x}_{k-i})\right)\right\|^{2}.$ (6)
@@
In the literature, many efforts have been made to improve the performance of AA. For example, in [14], the authors propose a systematic way to dynamically alternate the window size $m_{1}$ by the multiplicative composite combination, i.e., applying AA($m_{1}$) in the outer loop and applying AA($m_{2}$) in the inner loop. In [15], the authors consider a nonstationary Anderson acceleration algorithm with optimized damping in each iteration by applying one extra inexpensive optimization, i.e, adding another linear combination of previous $\{\mathbf{x}_{k-i}\}_{i=0}^{m_{k}}$ in (3). Rather than applying AA at every step, the authors in [31] employ AA at periodic intervals within the Jacobi iteration, named alternating Anderson-Jacobi method (AAR), which outperforms the GMRES method of problems considered. Subsequently, the alternating Anderson-Richardson method is extended to incorporate preconditioning, address efficient parallel implementation, and provide serial MATLAB and parallel C/C++ implementations [35]. The convergence analysis of alternating Anderson-Richardson method is presented in [27], along with a discussion of its connection to GMRES. We note that [27] uses the framework of Algorithm 2 to define AAR.
@@
In aAA(mm)[ss]–FP[$t$] when $t=0$, we recover AA($m$), for any given $s$. When $m=s=t=1$, aAA(mm)[ss]–FP[$t$] is restarted AA(1), which has been studied in [9, 26]. We note that when $s=1$, our aAA(mm)[ss]–FP[$t$] is not equivalent to the AAR presented in [27, Algorithm 1]. In fact, in AAR, if we set $p=2$ (the period) and $m=\infty$, the full AAR gives
@@
3 Convergence analysis
@@
where $\mathbf{A}\in\mathbb{R}^{n\times n}$ and $\mathbf{x}^{*}$ is the exact solution. We consider the Richardson iteration as the fixed-point iteration for solving (10), which is given by
@@
In the literature, it has been shown [37, 30] that under certain conditions the iterates, $\mathbf{x}_{k}$, of AA($\infty$) can be recovered from GMRES iterate $\mathbf{x}_{k-1}^{G}$, i.e., $\mathbf{x}_{k}=q(\mathbf{x}_{k-1}^{G})$. In [27], the author establishes a connection between full AAR and GMRES. We present a similar result for aAA($\infty$)[11]–FP[$t$], which can be derived following [27, Theorem 1].
@@
$\mathbf{x}_{k+1}=(1+\sum_{i=1}^{k}\gamma_{i})\mathbf{M}\mathbf{x}_{k}-\sum_{i=1}^{k}\gamma_{i}M\mathbf{x}_{k-i}+\mathbf{b}=\mathbf{M}\left((1+\sum_{i=1}^{k}\gamma_{i})\mathbf{x}_{k}-\sum_{i=1}^{k}\gamma_{i}\mathbf{x}_{k-i}\right)+\mathbf{b}=q(\mathbf{\hat{x}}_{k}),$
@@
Following the idea in [27, Theorem 1], we have $\mathbf{\hat{x}}_{k}=\mathbf{x}_{k}^{G}$, where k=jp1k=jp-1. Thus, we obtain the desired result. ∎
@@
From [23], we know that the least-squares problem defined in (4) is minimizing $\|\mathbf{M}^{-1}\mathbf{r}_{k+1}\|$, provided that $\mathbf{M}$ is invertible, that is,
$\|\mathbf{M}^{-1}\mathbf{r}_{k+1}\|=\min_{\gamma_{i}^{(k)}}\|r(\mathbf{x}_{k})+\sum_{i=1}^{m_{k}}\gamma^{(k)}_{i}\left(r(\mathbf{x}_{k})-r(\mathbf{x}_{k-i})\right)\|^{2},$
and
$\|\mathbf{M}^{-1}\mathbf{r}_{k+1}\|\leq\|\mathbf{M}^{-1}r_{k}\|.$ (13)
@@
According to the definition of aAA(mm)[ss]–FP[$t$], we know that its iterate $\mathbf{x}_{k+1}$ is either generated from AA($m$) or the fixed-point iteration. If $\mathbf{x}_{k+1}$ is obtained from AA($m$), then from [36, Theorem 2.1], we know $\|\mathbf{r}_{k+1}\|\leq c\|\mathbf{r}_{k}\|$. If $\mathbf{x}_{k+1}$ is obtained by applying one fixed-point iteration given by (11), then we have $\mathbf{r}_{k+1}=\mathbf{M}\mathbf{r}_{k}$. It follows that $\|\mathbf{r}_{k+1}\|\leq c\|\mathbf{r}_{k}\|$. As a result, $\|\mathbf{r}_{k+1}\|\leq c\|\mathbf{r}_{k}\|\leq c^{k}\|\mathbf{r}_{0}\|$ for all $k$. Since $c<1$, $\|\mathbf{r}_{k+1}\|\rightarrow 0$. ∎
@@
$\mathbf{X}_{k}=\begin{bmatrix}\mathbf{e}_{k-m_{k}+1}-\mathbf{e}_{k-m_{k}}&\mathbf{e}_{k-m_{k}+2}-\mathbf{e}_{k-m_{k}+1},&\cdots,&\mathbf{e}_{k}-\mathbf{e}_{k-1}\end{bmatrix}$.
@@
$\mathbf{S}_{k}=(\mathbf{D}_{k}+\mathbf{X}_{k})(\mathbf{D}_{k}^{T}\mathbf{D}_{k})^{-1}\mathbf{D}_{k}^{T}.$
@@
In [40], the authors present a one-step analysis of AA when $\mathbf{M}$ is symmetric, that is, comparing the solution error $\mathbf{e}_{k+1}$ after one step of AA, where $\{\mathbf{x}_{j}=q(\mathbf{x}_{j-1})\}_{j=1}^{k}$, to the solution error after $(k+1)$ fixed-point iterations. This idea can be adapted to perform the convergence analysis of our aAA(mm)[ss]–FP[$t$] method. We first extend the results of [40, Lemma 3.1], originally formulated for the symmetric case, to the more general scenario, where $\mathbf{M}$ is diagonalizable and invertible.
@@
Here, $\mathbf{\tilde{E}}=\mathbf{\Pi}(\mathbf{I}-\mathbf{K}(\mathbf{K}^{T}\mathbf{W}^{T}\mathbf{W}\mathbf{K})^{-1}\mathbf{K}^{T}\mathbf{W}^{T}\mathbf{W})\mathbf{\Pi}^{-1}$ with $\mathbf{\Pi}=\mathbf{\Lambda}(\mathbf{\Lambda}-\mathbf{I})^{-1}$ and $\mathbf{K}=\mathcal{K}_{m_{k}}(\mathbf{\Lambda},\mathbf{H}\mathbf{W}^{-1}\mathbf{e}_{0})$, where $\mathbf{H}=(\mathbf{\Lambda}-\mathbf{I})^{2}\mathbf{\Lambda}^{k-m_{k}}$, and $\mathbf{R}=\mathbf{I}-\mathbf{W}\mathbf{K}(\mathbf{K}^{T}\mathbf{W}^{T}\mathbf{W}\mathbf{K})^{-1}\mathbf{K}^{T}\mathbf{W}^{T}$.
@@
It can be shown that
$\mathbf{X}_{k}=\mathbf{W}(\mathbf{\Lambda}-\mathbf{I})^{-1}\mathbf{K},\quad\mathbf{D}_{k}=\mathbf{WK}.$
Moreover,
@@
$\mathbf{I}-(\mathbf{I}-\mathbf{S}_{k})\mathbf{A}$ $=\mathbf{M}+\mathbf{S}_{k}(\mathbf{I}-\mathbf{M})$
@@
$=\mathbf{W\Lambda}\mathbf{W}^{-1}+\mathbf{W}\mathbf{\Lambda K}(\mathbf{K}^{T}\mathbf{W}^{T}\mathbf{WK})^{-1}\mathbf{K}^{T}\mathbf{W}^{T}$
@@
We note that when $\mathbf{M}$ is symmetric, $\mathbf{W}$ is a unitary matrix, that is, $\mathbf{W}^{T}\mathbf{W}=\mathbf{I}$. Consequently, the relationship (15) coincides with the result given in [40, Lemma 3.1]. Next, we provide an estimate for $\|\mathbf{r}_{k+1}\|$ based on the above result.
@@
From Lemma 1, we have $\|\mathbf{r}_{k+1}\|\leq\|\mathbf{M}\|\|\mathbf{R}\mathbf{r}_{k}\|$. We then estimate $\|\mathbf{R}\mathbf{r}_{k}\|$.
@@
$=\min_{p\in\mathbb{P}_{m_{k}-1}}\|\mathbf{r}_{k}-p(\mathbf{M})\mathbf{WH}\mathbf{W}^{-1}\mathbf{e}_{0}\|.$
However, $\mathbf{W}^{-1}\mathbf{M}\mathbf{e}_{k}=\mathbf{W}^{-1}\mathbf{M}^{k+1}\mathbf{e}_{0}=\mathbf{\Lambda}^{k+1}\mathbf{W}^{-1}\mathbf{e}_{0}$. Thus,
@@
$=-\mathbf{W}(\mathbf{I}-\mathbf{\Lambda})\mathbf{\Lambda}^{-m_{k}}\mathbf{W}^{-1}\mathbf{r}_{k}.$

### v2

@@ 1 Introduction
Anderson acceleration (AA) or Anderson mixing [2] was first proposed by Donald G. Anderson in 1965 to improve the performance of fixed-point iterations in solving systems of nonlinear equations. Later, AA is applied to solve both linear and nonlinear problems. The main idea of AA is to use previous iterates to generate a new approximation. Overtime, AA has been widely used in many areas, such as fluid flow [9, 32], optimization problems [40, 7, 31, 20], reinforcement learning [35, 36, 41, 23], and so on [1, 25, 26]. Under certain conditions, when AA using all previous iteration information in combination with Richardson iteration is applied to solving linear systems, its iterates can be recovered from the GMRES method; see [39, 33]. Recently, there have been many studies on the convergence analysis of AA [33, 39, 38, 19, 16, 17, 8, 12] and its development [37, 34, 15, 14]. A recent review on AA can be found in [3].
@@
In practice, we consider AA with finite window size, $m$, referred to as AA($m$). In AA($m$), each update is a linear combination of the underlying fixed-point iteration acting on previous $m+1$ iterates. The combination coefficients change at each iteration and are obtained by solving a least-squares problem. To reduce computational cost, one may consider a different combination of AA and fixed-point iteration. In [34], AA is employed at periodic intervals within Jacobi iteration. Later, [37] applies this strategy to Richardson iteration, where a preconditioner is considered. [30] further provides a convergence analysis of the alternating Anderson with Jacobi iteration when the underlying fixed-point iteration is contractive and makes a connection with GMRES for full alternating Anderson, i.e., using all previous iteration information. A convergence analysis of alternating Anderson-Picard, in which an AA($m$) step is applied after $m$ steps of Picard iterations, for nonlinear problems, is provided in [21]. In addition, the authors in [21] establish the equivalence between the alternating Anderson-Picard and a multisecant-GMRES method that employs GMRES to solve a multisecant linear system at each iteration, and demonstrate the efficiency and robustness of the alternating Anderson-Picard. In [20, Algorithm 7], the author considers a Generalized Alternate Anderson-Picard (GAAP) method, which allows flexible configurations of the Picard step $t$ and the Anderson acceleration history $m$. In this work, we extend the alternating strategy to a general mixing scheme, where we consider a periodic pattern comprising $t$ steps of fixed-point iteration interleaved with $s$ steps of AA($m$). We refer to this approach as aAA(mm)[ss]–FP[$t$] or simply aAA–FP. Here, $m,s,t$ are integers. Note that the GAAP method is aAA(mm)[1]–FP[$t$]. In [4], it was found that taking $m\geq 2t$ generally improves performance of aAA(mm)[1]–FP[$t$] for the problems considered there. Our aAA–FP method provides us with more flexibility for general models compared to the GAAP method. In this work, we investigate the selection of parameters $m$, $s$, and $t$ to improve the performance of aAA–FP.
@@
  • Third, to highlight its wide-ranging utility, our method is employed to enhance the efficiency of Jacobi iteration, Gauss–Seidel iteration, Picard iteration, gradient descent, and the alternating direction method of multipliers for solving partial differential equations and nonlinear optimization problems. We provide numerical evidence of the efficiency of aAA–FP in terms of both iteration count and CPU time, compared to the standalone fixed-point iteration, GMRES, and aAA(mm)[11]–FP[$t$]. The results show that, for suitable choices of the parameters, $m,s,t$, aAA(mm)[ss]–FP[$t$] is competitive with, or outperforms, the other methods.
Regarding the numerical results, we conduct a wide range of test problems, including nonsymmetric linear systems, Navier–Stokes equations, and nonlinear, nonsmooth optimization problems. For the highly ill-conditioned, nonsymmetric linear systems, we consider aAA–FP, GMRES, windowed AA, and AA($\infty$) that are applied to accelerate the noncontractive Jacobi and contractive Gauss–Seidel iterations. Our numerical results show that aAA–FP with Gauss-Seidel outperforms the Jacobi-based aAA–FP version. For the Navier–Stokes problem, we accelerate the Picard iteration by aAA–FP and AA. We see that aAA–FP significantly speeds up the performance of the Picard iteration. Moreover, aAA–FP can be better than AA($\infty$). For the nonlinear optimization problems studied here, we consider the Alternating Direction Method of Multipliers (ADMM) as a fixed-point iteration. Our study demonstrates that aAA–FP and AA accelerate ADMM by approximately an order of magnitude, resulting in a reduction by a factor of 10 in both CPU time and iteration count. We also note that for some careful choices of $s$ and $t$, aAA(mm)[ss]–FP[$t$] outperforms AA($m$) and is competitive as AA($\infty$). Finally, we accelerate the gradient descend (GD) in the solution of the regularized logistic regression. Also for this problem, we observe that AA and aAA–FP can drastically speed-up the solution process by an order of magnitude, with aAA–FP being able to outperform AA for suitable choice of the parameters. In conclusion, our aAA(mm)[ss]–FP[$t$] method has the potential to solve challenging linear and nonlinear problems for which GMRES is not applicable in the nonlinear case, and for which the underlying fixed-point iteration may converge very slowly or even diverge.
@@ 2 Generalized alternating Anderson acceleration
$\min_{\left(\gamma_{1}^{(k)},\cdots,\gamma_{m_{k}}^{(k)}\right)}\left\|r(\mathbf{x}_{k})+\sum_{i=1}^{m_{k}}\gamma_{i}^{(k)}\left(r(\mathbf{x}_{k})-r(\mathbf{x}_{k-i})\right)\right\|^{2},$ (4)
where $r(\mathbf{x})=q(\mathbf{x})-\mathbf{x}$.
@@
$\min_{\left(\tau_{1}^{(k)},\cdots,\tau_{k}^{(k)}\right)}\left\|r(\mathbf{x}_{k})+\sum_{i=1}^{m_{k}}\tau^{(k)}_{i}\left(r(\mathbf{x}_{k-i+1})-r(\mathbf{x}_{k-i})\right)\right\|^{2},$ (6)
where $r(\mathbf{x})=q(\mathbf{x})-\mathbf{x}$.
@@
In the literature, many efforts have been made to improve the performance of AA. For example, in [14], the authors propose a systematic way to dynamically alternate the window size $m_{1}$ by the multiplicative composite combination, i.e., applying AA($m_{1}$) in the outer loop and applying AA($m_{2}$) in the inner loop. In [15], the authors consider a nonstationary Anderson acceleration algorithm with optimized damping in each iteration by applying one extra inexpensive optimization, i.e, adding another linear combination of previous $\{\mathbf{x}_{k-i}\}_{i=0}^{m_{k}}$ in (3). Rather than applying AA at every step, the authors in [34] employ AA at periodic intervals within the Jacobi iteration, named alternating Anderson-Jacobi method (AAJ), which outperforms the GMRES method of problems considered. Subsequently, the alternating Anderson-Richardson (AAR) method is extended to incorporate preconditioning, address efficient parallel implementation, and provide serial MATLAB and parallel C/C++ implementations [37]. The convergence analysis of AAR method is presented in [30], along with a discussion of its connection to GMRES. We note that [30] uses the framework of Algorithm 2 to define AAR.
@@
In aAA(mm)[ss]–FP[$t$] when $t=0$, we recover AA($m$), for any given $s$. When $m=s=t=1$, aAA(mm)[ss]–FP[$t$] is restarted AA(1), which has been studied in [9, 27]. We note that when $s=1$, our aAA(mm)[ss]–FP[$t$] is not equivalent to the AAR presented in [30, Algorithm 1]. In fact, the AAR method uses $s=1$, $p=s+t$, defining the periodicity, and sets $x_{1}=q(x_{0})$. If we set $p=2$ and $m=\infty$, the full AAR gives
@@
3 Convergence analysis for the linear case
@@
where $\mathbf{A}\in\mathbb{R}^{n\times n}$ and $\mathbf{x}^{*}\in\mathbb{R}^{n}$ is the exact solution. We consider the Richardson iteration as the fixed-point iteration for solving (10), which is given by
@@
In the literature, it has been shown [39, 33] that under certain conditions the iterates, $\mathbf{x}_{k}$, of AA($\infty$) can be recovered from GMRES iterate $\mathbf{x}_{k-1}^{G}$, i.e., $\mathbf{x}_{k}=q(\mathbf{x}_{k-1}^{G})$. In [30], the author establishes a connection between full AAR and GMRES. We present a similar result for aAA($\infty$)[11]–FP[$t$], which can be derived following [30, Theorem 1].
@@
$\mathbf{x}_{k+1}=(1+\sum_{i=1}^{k}\gamma_{i})\mathbf{M}\mathbf{x}_{k}-\sum_{i=1}^{k}\gamma_{i}M\mathbf{x}_{k-i}+\mathbf{b}=\mathbf{M}((1+\sum_{i=1}^{k}\gamma_{i})\mathbf{x}_{k}-\sum_{i=1}^{k}\gamma_{i}\mathbf{x}_{k-i})+\mathbf{b}=q(\mathbf{\hat{x}}_{k}),$
@@
Following the idea in [30, Theorem 1], we have $\mathbf{\hat{x}}_{k}=\mathbf{x}_{k}^{G}$, where k=jp1k=jp-1. Thus, we obtain the desired result. ∎
@@
For the AA($m$) method, [24] has shown that the least-squares problem defined in (4) is minimizing $\|\mathbf{M}^{-1}\mathbf{r}_{k+1}\|$, provided that $\mathbf{M}$ is invertible, that is,
$\min_{\left(\gamma_{1}^{(k)},\cdots,\gamma_{k}^{(k)}\right)}\|\mathbf{M}^{-1}\mathbf{r}_{k+1}\|=\min_{\left(\gamma_{1}^{(k)},\cdots,\gamma_{k}^{(k)}\right)}\|r(\mathbf{x}_{k})+\sum_{i=1}^{m_{k}}\gamma^{(k)}_{i}\left(r(\mathbf{x}_{k})-r(\mathbf{x}_{k-i})\right)\|^{2}.$
Moreover, [24] has established that
$\|\mathbf{M}^{-1}\mathbf{r}_{k+1}\|\leq\|r_{k}\|.$ (13)
@@
According to the definition of aAA(mm)[ss]–FP[$t$], we know that its iterate $\mathbf{x}_{k+1}$ is either generated from AA($m$) or the fixed-point iteration. If $\mathbf{x}_{k+1}$ is obtained from AA($m$), then from [38, Theorem 2.1], we know $\|\mathbf{r}_{k+1}\|\leq c\|\mathbf{r}_{k}\|$. If $\mathbf{x}_{k+1}$ is obtained by applying one fixed-point iteration given by (11), then we have $\mathbf{r}_{k+1}=\mathbf{M}\mathbf{r}_{k}$. It follows that $\|\mathbf{r}_{k+1}\|\leq c\|\mathbf{r}_{k}\|$. As a result, $\|\mathbf{r}_{k+1}\|\leq c\|\mathbf{r}_{k}\|\leq c^{k+1}\|\mathbf{r}_{0}\|$ for all $k$. Since $c<1$, $\|\mathbf{r}_{k+1}\|\rightarrow 0$. ∎
@@
$\mathbf{X}_{k}=\begin{bmatrix}\mathbf{e}_{k-m_{k}+1}-\mathbf{e}_{k-m_{k}}&\mathbf{e}_{k-m_{k}+2}-\mathbf{e}_{k-m_{k}+1},&\cdots,&\mathbf{e}_{k}-\mathbf{e}_{k-1}\end{bmatrix}$. (14)
@@
$\mathbf{S}_{k}=(\mathbf{D}_{k}+\mathbf{X}_{k})(\mathbf{D}_{k}^{T}\mathbf{D}_{k})^{-1}\mathbf{D}_{k}^{T}.$ (15)
@@
In [42], the authors present a one-step analysis of AA when $\mathbf{M}$ is symmetric, that is, comparing the solution error $\mathbf{e}_{k+1}$ after one step of AA, where $\{\mathbf{x}_{j}=q(\mathbf{x}_{j-1})\}_{j=1}^{k}$, to the solution error after $(k+1)$ fixed-point iterations. This idea can be adapted to perform the convergence analysis of our aAA(mm)[ss]–FP[$t$] method. We first extend the results of [42, Lemma 3.1], originally formulated for the symmetric case, to the more general scenario, where $\mathbf{M}$ is diagonalizable and invertible.
@@
Here, $\mathbf{\tilde{E}}=\mathbf{\Pi}(\mathbf{I}-\mathbf{K}(\mathbf{K}^{T}\mathbf{W}^{T}\mathbf{W}\mathbf{K})^{-1}\mathbf{K}^{T}\mathbf{W}^{T}\mathbf{W})\mathbf{\Pi}^{-1}$ with $\mathbf{\Pi}=\mathbf{\Lambda}(\mathbf{\Lambda}-\mathbf{I})^{-1}$ and $\mathbf{K}=\mathcal{K}_{m_{k}}(\mathbf{\Lambda},\mathbf{H}\mathbf{W}^{-1}\mathbf{e}_{0})$, where $\mathbf{H}=(\mathbf{\Lambda}-\mathbf{I})^{2}\mathbf{\Lambda}^{k-m_{k}}$, and
$\mathbf{R}=\mathbf{I}-\mathbf{W}\mathbf{K}(\mathbf{K}^{T}\mathbf{W}^{T}\mathbf{W}\mathbf{K})^{-1}\mathbf{K}^{T}\mathbf{W}^{T}.$ (18)
@@
We first show that
$\mathbf{X}_{k}=\mathbf{W}(\mathbf{\Lambda}-\mathbf{I})^{-1}\mathbf{K}\quad\text{and}\quad\mathbf{D}_{k}=\mathbf{WK}.$
For $1\leq j\leq k$, $\mathbf{x}_{j}=\mathbf{M}\mathbf{x}_{j-1}+b$, we have $\mathbf{e}_{j}=\mathbf{M}^{j}\mathbf{e}_{0}$. It follows that
$\mathbf{e}_{j}-\mathbf{e}_{j-1}=(\mathbf{W\Lambda}^{j}\mathbf{W}^{-1}-\mathbf{W\Lambda}^{j-1}\mathbf{W}^{-1})\mathbf{e}_{0}=\mathbf{W(\Lambda-I)\Lambda}^{j-1}\mathbf{W}^{-1}\mathbf{e}_{0}.$
From (14), we have
$\mathbf{X}_{k}$ $=\mathbf{W(\Lambda-I)}\mathbf{\Lambda}^{k-m_{k}}[\mathbf{W}^{-1}\mathbf{e}_{0},\mathbf{\Lambda W}^{-1}\mathbf{e}_{0},\cdots,\mathbf{\Lambda}^{m_{k}-1}\mathbf{W}^{-1}\mathbf{e}_{0}]$
$=\mathbf{W(\Lambda-I)}^{-1}(\mathbf{\Lambda-I})^{2}\mathbf{\Lambda}^{k-m_{k}}[\mathbf{W}^{-1}\mathbf{e}_{0},\mathbf{\Lambda W}^{-1}\mathbf{e}_{0},\cdots,\mathbf{\Lambda}^{m_{k}-1}\mathbf{W}^{-1}\mathbf{e}_{0}]$
$=\mathbf{W(\Lambda-I)}^{-1}[\mathbf{HW}^{-1}\mathbf{e}_{0},\mathbf{\Lambda HW}^{-1}\mathbf{e}_{0},\cdots,\mathbf{\Lambda}^{m_{k}-1}\mathbf{HW}^{-1}\mathbf{e}_{0}]$
$=\mathbf{W(\Lambda-I)}^{-1}\mathcal{K}_{m_{k}}(\mathbf{\Lambda},\mathbf{H}\mathbf{W}^{-1}\mathbf{e}_{0}),$
$=\mathbf{W(\Lambda-I)}^{-1}\mathbf{K},$
which is the desired result.
Since $\mathbf{D}_{k}=-\mathbf{A}\mathbf{X}_{k}$ and $\mathbf{A}=\mathbf{I}-\mathbf{M}$, we have
$\mathbf{D}_{k}=-(\mathbf{I}-\mathbf{W\Lambda W}^{-1})\mathbf{W}(\mathbf{\Lambda}-\mathbf{I})^{-1}\mathbf{K}=\mathbf{WK}.$
Moreover, using (15) leads to
@@
$\mathbf{I}-(\mathbf{I}-\mathbf{S}_{k})\mathbf{A}$
$=$ $\mathbf{M}+\mathbf{S}_{k}(\mathbf{I}-\mathbf{M})$
@@
$=\mathbf{W\Lambda}\mathbf{W}^{-1}-\mathbf{W}\mathbf{\Lambda K}(\mathbf{K}^{T}\mathbf{W}^{T}\mathbf{WK})^{-1}\mathbf{K}^{T}\mathbf{W}^{T}$
@@
We note that when $\mathbf{M}$ is symmetric, $\mathbf{W}$ is a unitary matrix, that is, $\mathbf{W}^{T}\mathbf{W}=\mathbf{I}$. Consequently, the relationship (17) coincides with the result given in [42, Lemma 3.1]. Next, we provide an estimate for $\|\mathbf{r}_{k+1}\|$ based on the above result.
@@
From Lemma 1, we have $\|\mathbf{r}_{k+1}\|\leq\|\mathbf{M}\|\|\mathbf{R}\mathbf{r}_{k}\|$. We then estimate $\|\mathbf{R}\mathbf{r}_{k}\|$. Using the definition of $\mathbf{R}$ in (18), we have
@@
$=\min_{p_{\ell}\in\mathbb{P}_{m_{k}-1}}\|\mathbf{r}_{k}-p_{\ell}(\mathbf{M})\mathbf{WH}\mathbf{W}^{-1}\mathbf{e}_{0}\|.$
However, for $1\leq j\leq k,\mathbf{x}_{j}=q(\mathbf{x}_{j-1})=\mathbf{M}\mathbf{x}_{j-1}+b$. It follows that $\mathbf{e}_{j}=\mathbf{M}\mathbf{x}_{j-1}+\mathbf{b}-\mathbf{x}^{*}=\mathbf{M}\mathbf{e}_{j-1}=\mathbf{M}^{j}\mathbf{e}_{0}$. Then, $\mathbf{W}^{-1}\mathbf{M}\mathbf{e}_{k}=\mathbf{W}^{-1}\mathbf{M}^{k+1}\mathbf{e}_{0}=\mathbf{\Lambda}^{k+1}\mathbf{W}^{-1}\mathbf{e}_{0}$. Thus,
@@
$=-\mathbf{W}(\mathbf{I}-\mathbf{\Lambda})\mathbf{\Lambda}^{-m_{k}}\mathbf{W}^{-1}\mathbf{r}_{k},$
where in the second-to-last equality we use the relation $\mathbf{r}_{k}=-\mathbf{A}\mathbf{e}_{k}=-(\mathbf{I-M})\mathbf{e}_{k}$.